if one piece of square tile measures 8 inches by 8 inches, what is the least number of tiles they will need to cover the patio floor if the square footage is 168sq.ft.
Answer by Alan3354(69443)

8 in = 2/3 ft
Area of tiles = (2/3)^2 = 4/9 sq ft
------------
168/(4/9) = 378 tiles